# A stochastic reconstruction theorem on rectangular increments with an application to a mixed hyperbolic SPDE

### v2

@@ 4.2 Products of processes and distributions
We now consider the case of the product (4.3) when $\zeta=\xi$ is white noise. Since $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$ regularises with “infinite” regularity the conditional expectation we allow ourselves to consider integrands $u\in C^{{\bm{\alpha}}}L_{m}$. In this case, the reconstruction coincides with the stochastic integral with respect to the Brownian sheet, see [Ker23, Section 7], [CW75] so we expect the reconstruction $f$ to have more properties than the ones given by the reconstruction theorem.
@@
Let $u\in C^{{\bm{\alpha}}}L_{m}$, ${\bm{\alpha}}>{\bm{0}}$ and $\xi$ be a white noise over +$\mathbb{R}^{d}_{+}$. Then the product germ $u\cdot\xi\in G^{-\frac{1}{2}{\bm{1}},-\frac{1}{2}{\bm{1}}+{\bm{\alpha}},\infty}L_{m}$ and for all test functions $\psi\in C^{r}_{c}((0,T)^{d})$, we have the identity
@@ 4.3 Primitives and integration
$\mathbf{y}\mapsto\oint_{\mathbf{s}}^{\mathbf{y}}f(d\mathbf{z})=:I_{\mathbf{s},\mathbf{x}}(f)$
@@
$\left\|f(\hat{P}_{\mathbf{m}}^{\chi}1_{[\mathbf{s},\mathbf{t}]})\right\|_{m}\leq\sum_{\mathbf{x}\in\Delta_{\mathbf{m}}}\left\|f(\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}})\right\|_{m}\left|\left\langle\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}},1_{(\mathbf{s},\mathbf{t}]}\right\rangle\right|\,.$
@@
Letting $\mathbf{k}\to\infty$, (4.11) then follows from $2^{-n_{i}}\approx\epsilon_{i}\left|t_{i}-s_{i}\right|$. The estimate (4.12) follows analogously. To see the identity (4.10), we simply observe that $P_{\mathbf{n}}1_{(\mathbf{s},\mathbf{t}]}$ is additive for each $\mathbf{n}\in\mathbb{N}^{d}$. ∎
Due to the presence of the variable ${\bm{\epsilon}}$ in Proposition 4.7, the interior primitive of $f$ does not suffice to define a proper integration over the whole cube $[0,T]^{d}$ when $\mathbf{s}$ and $\mathbf{t}$ approach the boundary. If only $\mathbf{s}$ or $\mathbf{t}$ approaches the boundary while the other one has some fixed distance to it, we can absorb this singularity via a dyadic expansion. Thanks to the additive property (4.10), this is an extension to the interior primitive.
@@
It follows that $Y\in C^{{\bm{\alpha}}+{\bm{1}},{\bm{\delta}}}L_{m}$. We call $Y$ the primitive, and we have
@@ 4.4 Existence and uniqueness
We are now ready to state the main theorem for the existence and uniqueness of a mild solution (4.2) on any finite time interval $[0,T]^{d}$ and an $m\geq 2$.
@@
Let $\xi$ be white noise over +$\mathbb{R}^{d}_{+}$ and $Z\in C^{{\bm{\beta}}}([0,T]^{d})$ for some $\frac{1}{2}{\bm{1}}<{\bm{\beta}}<{\bm{1}}$. Let also be ${\bm{\delta}}={\bm{\beta}}-\frac{1}{2}{\bm{1}}$ and ${\bm{\alpha}}$ satisfying ${\bm{\delta}}\leq{\bm{\alpha}}<\frac{1}{2}{\bm{1}}$ and ${\bm{\delta}}+{\bm{\beta}}+{\bm{\alpha}}>{\bm{1}}$. Under the assumption $\sigma\in\mathcal{C}^{1}(\mathbb{R})$ and $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}([0,T]^{d})$ for any initial datum $v\colon\partial_{{\bm{0}}}[0,T]^{d}\to\mathbb{R}$ such that $\mathcal{I}(v)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ there exists a unique solution $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ to the equation (4.2)
@@ 5 Connection to stochastic multiparameter sewing
In dimension $d=1$, sewing can be seen as a special case of the reconstruction theorem [BZ22]. This observation also extends to the stochastic case [Ker23], but in general no longer holds in dimension $d>1$. Indeed, Harang’s multiparameter sewing lemma [Har21] and the reconstruction theorem described in [CZ20] are not connected in similarly. We claim that our formulation using rectangular increment is the missing link between the two results. More precisely, this section shows that Theorem 3.12 extends the stochastic multiparameter sewing lemma from [BHK23]. Let us start by recalling the sewing lemma in question. Throughout this section, we fix a commuting filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$.
A germ in the world of sewing is a $2d$-parameter process given by a map $\Xi:\Delta_{T}\times\Omega\to\mathbb{R}$, where
@@
$\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}=L_{m}-\lim_{\left|\mathcal{P}\right|\to 0}\sum_{[\mathbf{u},\mathbf{v}]\in\mathcal{P}^{\theta}}\Xi_{\mathbf{u},\mathbf{v}}\,.$

### v3

@@ 4.2 Products of processes and distributions
We now consider the case of the product (4.3) when $\zeta=\xi$ is white noise. Since $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$ regularise with “infinite” regularity the conditional expectation from the filtration generated by the Brownian sheet of $\xi$, we allow ourselves to consider integrands $u\in C^{{\bm{\alpha}}}L_{m}$. In this case, the reconstruction coincides with the stochastic integral with respect to the Brownian sheet, see [Ker24, Section 7], [CW75] so we expect the reconstruction $f$ to have more properties than the ones given by the reconstruction theorem.
@@
Let $\xi$ be a white noise over +$\mathbb{R}^{d}_{+}$ and $u\in C^{{\bm{\alpha}}}L_{m}$, ${\bm{\alpha}}>{\bm{0}}$ is adapted to the filtration generated by the Brownian sheet of $\xi$. Then the product germ $u\cdot\xi\in G^{-\frac{1}{2}{\bm{1}},-\frac{1}{2}{\bm{1}}+{\bm{\alpha}},\infty}L_{m}$ and for all test functions $\psi\in C^{r}_{c}((0,T)^{d})$, we have the identity
@@ 4.3 Primitives and integration
$\mathbf{y}\mapsto\oint_{\mathbf{s}}^{\mathbf{y}}f(d\mathbf{z})=:I_{\mathbf{s},\mathbf{y}}(f)$
@@
$\left\|f(\hat{P}_{\mathbf{m}}^{\chi}1_{(\mathbf{s},\mathbf{t}]})\right\|_{m}\leq\sum_{\mathbf{x}\in\Delta_{\mathbf{m}}}\left\|f(\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}})\right\|_{m}\left|\left\langle\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}},1_{(\mathbf{s},\mathbf{t}]}\right\rangle\right|\,.$
@@
Since ${\bm{\alpha}}>-{\bm{1}}$, one has
$\sum_{\mathbf{m}_{\chi}\geq\mathbf{n}_{\chi}}\left\|f(\hat{P}_{\mathbf{m}}^{\chi}1_{(\mathbf{s},\mathbf{t}]})\right\|_{m}<\infty\,.$
@@
By letting $\mathbf{k}\to\infty$, the first inequality (4.11) then from $2^{-n_{i}}\approx\epsilon_{i}\left|t_{i}-s_{i}\right|$. The estimate (4.12) follows analogously. To see the identity (4.10), we simply observe that $P_{\mathbf{n}}1_{(\mathbf{s},\mathbf{t}]}$ is additive for each $\mathbf{n}\in\mathbb{N}^{d}$. ∎
At this point, the interior integral $I_{\mathbf{s},\mathbf{y}}(f)$ might depend on the choice of our wavelet basis $\Phi$. We show now that the limit is unique by showing that its mixed distributional derivative is $f$. That, together with the fact that $I$ vanishes on the boundary by construction, that is, $I_{\mathbf{s},\pi^{i}_{s_{i}}\mathbf{x}}=0$ for all $\mathbf{s},\mathbf{x}\in(0,T^{d})$ and $i\in[d]$, uniquely characterizes $(\mathbf{s},\mathbf{y})\mapsto I_{\mathbf{s},\mathbf{y}}$.
Lemma 4.10.
For each ${\bm{0}}<\mathbf{s}\leq\mathbf{y}<\mathbf{T}$ let $I_{\mathbf{s},\mathbf{y}}$ be as above. Further let $\psi\in C_{c}^{rd}$ with support in $(\mathbf{s},\mathbf{T})$. Then it holds that
$(-1)^{d}\left\langle I_{\mathbf{s},\cdot},\partial^{\bm{1}}\psi\right\rangle=f(\psi)\,.$
Proof.
Direct computation gives us for any $\mathbf{n}\in\mathbb{N}^{d}$
[…]
$=\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}f(\phi_{\mathbf{x}}^{\mathbf{n}})\left\langle\phi_{\mathbf{x}}^{\mathbf{n}},\psi\right\rangle=f(P_{\mathbf{n}}\psi)\,.$
[…]
Let us analyse the convergence of the left-hand side. Using Minkowski inequality, (4.13) and (4.14), we get
[…]
$\lesssim\left\|\psi\right\|_{C^{rd}}\int_{\bm{0}}^{\mathbf{T}}\sum_{\emptyset\neq\chi\subset[d]}\sum_{\begin{subarray}{c}\mathbf{m}_{\chi}\geq\mathbf{n}_{\chi}\\ m_{\chi^{c}}=\mathbf{n}_{\chi^{c}}\end{subarray}}\left\|f(\hat{P}^{\chi}_{\mathbf{m}}1_{(\mathbf{s},\mathbf{z}]})\right\|_{m}d\mathbf{z}$
$\lesssim\left\|\psi\right\|_{C^{rd}}\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\sum_{\emptyset\neq\chi\subset[d]}2^{-\mathbf{n}*{\bm{\alpha}}}2^{-\mathbf{n}}_{\chi}\mathbf{T}^{2\cdot{\bm{1}}}\xrightarrow{\mathbf{n}\to\infty}0\,.$
It follows that
$(-1)^{d}\left\langle f(P_{\mathbf{n}}1_{(\mathbf{s},\cdot]}),\partial^{\bm{1}}\psi\right\rangle\xrightarrow{\mathbf{n}\to\infty}(-1)^{d}\left\langle I_{\mathbf{s},\cdot},\partial^{\bm{1}}\psi\right\rangle\,,$
finishing the proof. ∎
Due to the presence of the variable ${\bm{\epsilon}}$ in Proposition 4.7, the interior primitive of $f$ does not suffice to define a proper integration over the whole cube $[0,T]^{d}$ when $\mathbf{s}$ and $\mathbf{t}$ approach the boundary. If only $\mathbf{s}$ or $\mathbf{t}$ approaches the boundary while the other one has some fixed distance to it, we can absorb this singularity via a dyadic expansion. Due to the additive property (4.10), this is an extension of the interior primitive.
@@
It follows that $Y\in C^{{\bm{\alpha}}+{\bm{1}},{\bm{\delta}}}L_{m}$. We call $Y$ the primitive of $f$, and we have
@@ 4.4 Existence and uniqueness
We are now ready to state the main theorem for the existence and uniqueness of (4.2) on any finite time interval $[0,T]^{d}$ and an $m\geq 2$.
@@
Let $\xi$ be white noise over +$\mathbb{R}^{d}_{+}$ and $Z\in C^{{\bm{\beta}}}([0,T]^{d})$ for some $\frac{1}{2}{\bm{1}}<{\bm{\beta}}<{\bm{1}}$. Let also be ${\bm{\delta}}={\bm{\beta}}-\frac{1}{2}{\bm{1}}$ and ${\bm{\alpha}}$ satisfying ${\bm{\delta}}\leq{\bm{\alpha}}<\frac{1}{2}{\bm{1}}$ and ${\bm{\delta}}+{\bm{\beta}}+{\bm{\alpha}}>{\bm{1}}$. Under the assumption $\sigma\in\mathcal{C}^{1}(\mathbb{R})$ and $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}([0,T]^{d})$ for any initial datum $v\colon\partial_{{\bm{0}}}[0,T]^{d}\to\mathbb{R}$ such that $I(v)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ there exists a unique solution $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ to the equation (4.2), with $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ defined with the filtration generated by the Brownian sheet of $\xi$.
@@ 5 Connection to stochastic multiparameter sewing
In dimension $d=1$, sewing can be seen as a special case of the reconstruction theorem [BZ22]. This observation also extends to the stochastic case [Ker24], but in general does not hold in dimension $d>1$. In fact, Harang’s multiparameter sewing lemma [Har21] and the reconstruction theorem described in [CZ20] are not connected in a similar way. We claim that our formulation using rectangular increments is the missing link between the two results. As in dimension $1$, Theorem 3.12 extends the stochastic multiparameter sewing lemma from [BHK23]. It also goes beyond it, as reconstruction is able to handle negative regularities. In this section, we show how one can derive the stochastic multiparameter sewing lemma from Theorem 3.12 and then present an example which lies beyond the scope of sewing, but can be solved by our reconstruction theorem.
5.1 Deriving sewing from reconstruction
Let us start by recalling the sewing lemma in question. Throughout this section, we fix a commuting filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$. A germ in the world of sewing is a $2d$-parameter process given by a map $\Xi:\Delta_{T}\times\Omega\to\mathbb{R}$, where
@@ 5.1 Deriving sewing from reconstruction
$\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}=\lim_{\left|\mathcal{P}\right|\to 0}\sum_{[\mathbf{u},\mathbf{v}]\in\mathcal{P}^{\theta}}\Xi_{\mathbf{u},\mathbf{v}}\,\quad\text{in}\;L_{m}.$